# FINALLY: fast and universal speech enhancement with studio-like quality

### v1

@@ C.4 Limitations
Appendix D Evaluation details
D.1 Metric computation
To compare models in their ability to correctly reconstruct the speech we use Phoneme Error Rate (PhER) and Word Error Rate (WER). Given the reference audio and audio, produced by the speech enhancement model, metrics are computed in the following way. First, Automatic Speech Recognition (ASR) model is applied to both audios, yielding phoneme representations. Then, the editing distance is computed over predicted phonemes (or words for WER) for the reference and the predicted audio. The distance is normalized to the length of the reference phoneme representation (or the length of the sentence for WER) and averaged over the validation set. We rely on https://huggingface.co/ for the ASR models and metric computations. For PhER we use mrrubino/wav2vec2-large-xlsr-53-l2-arctic-phoneme model, for WER we use jonatasgrosman/wav2vec2-large-xlsr-53-english. These are Wav2Vec2-based (Baevski et al., 2020) models for ASR.
D.2 Ablation details
For the ablation study, we utilized a smaller version of our original model to expedite evaluations, with the reduced model containing approximately 22 million parameters. The dataset used for this ablation was provided by (Nagrani et al., 2017). The ablation process was carried out in several stages. Initially, we compared the performance of our LMOS loss with L1Spec loss (Kong et al., 2020), MS-STFT loss (Defossez et al., 2020), and RecLoss (Défossez et al., 2023) during the pretraining phase. Each loss function was used independently to train the model in a regressive manner for 500,000 steps with a batch size of 12, utilizing two Nvidia P40 GPUs. Our findings indicated that LMOS consistently outperformed the other losses. Detailed results are presented in Table 11. It’s important to note that MOS evaluations were not conducted during the first stage, as the enhanced audio typically displayed artifacts following this stage of training.
@@ D.2 Ablation details
Appendix E Subjective evaluation
@@ 23. Answer: [No]
  24. Justification: We do not open the code for training of the model due to our organization policy, however, we use only publicly available data and most of the source codes are available as stated in Appendix C. Therefore, we expect that the work could be reproduced with moderate efforts.
@@ 28. Answer: [Yes]
  29. Justification: Please see the Appendix C
@@ 38. Answer: [Yes]
  39. Justification: lease see the Appendix C
@@ 58. Answer: [Yes]
  59. Justification: -
  60. Guidelines: Please the Section 5.
@@ 68. Answer: [Yes]
  69. Justification: Please see the Appendix E

### v2

@@ D.4 Limitations
Appendix E Evaluation details
E.1 Metric computation
To compare models in their ability to correctly reconstruct the speech, we use Phoneme Error Rate (PhER) and Word Error Rate (WER). Given the reference audio and audio, produced by the speech enhancement model, metrics are computed in the following way. First, Automatic Speech Recognition (ASR) model is applied to both audios, yielding phoneme representations. Then, the editing distance is computed over predicted phonemes (or words for WER) for the reference and the predicted audio. The distance is normalized to the length of the reference phoneme representation (or the length of the sentence for WER) and averaged over the validation set. We rely on https://huggingface.co/ for the ASR models and metric computations. For PhER we use mrrubino/wav2vec2-large-xlsr-53-l2-arctic-phoneme model, for WER we use jonatasgrosman/wav2vec2-large-xlsr-53-english. These are Wav2Vec2-based (Baevski et al., 2020) models for ASR.
E.2 Ablation details
For the ablation study, we utilized a smaller version of our original model to expedite evaluations, with the reduced model containing approximately 22 million parameters. The dataset used for this ablation was provided by (Nagrani et al., 2017). The ablation process was carried out in several stages. Initially, we compared the performance of our LMOS loss with L1Spec loss (Kong et al., 2020), MS-STFT loss (Defossez et al., 2020), and RecLoss (Défossez et al., 2023) during the pretraining phase. Each loss function was used independently to train the model in a regressive manner for 500,000 steps with a batch size of 12, utilizing two Nvidia P40 GPUs. Our findings indicated that LMOS consistently outperformed the other losses. Detailed results are presented in Table 11. It’s important to note that MOS evaluations were not conducted during the first stage, as the enhanced audio typically displayed artefacts following this stage of training.
@@ E.2 Ablation details
Appendix F Subjective evaluation
@@ 23. Answer: [No]
  24. Justification: We do not open the code for training of the model due to our organization policy, however, we use only publicly available data and most of the source codes are available as stated in Appendix D. Therefore, we expect that the work could be reproduced with moderate efforts.
@@ 28. Answer: [Yes]
  29. Justification: Please see the Appendix D
@@ 38. Answer: [Yes]
  39. Justification: please see the Appendix D
@@ 58. Answer: [Yes]
  59. Justification: Please the Section 5.
  60. Guidelines:
@@ 68. Answer: [Yes]
  69. Justification: Please see the Appendix F